OpenStudy (anonymous):

factor the perfect square trinomial completely. z^4-14z^2+49

OpenStudy (anonymous):

do you know how to factor trinomials?

OpenStudy (anonymous):

in quadratic form?

OpenStudy (anonymous):

 not really o.o

OpenStudy (anonymous):

well you know a perfect square trinomial factors into two binomials: \[(x + y)(x + y) \] the signs may vary. do you know this much?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay then let's get started disregard the variables for now, and find two numbers who's product is +49 and who's sum is -14

OpenStudy (anonymous):

-7 and -7

OpenStudy (anonymous):

okay you know that your constants in the binomial are going to be -7 and -7 so your equation will look something like this (x-7)(x-7) right? now all we need to do is find out what your x's are

OpenStudy (anonymous):

the leading variable is \[z^4\] and the middle variable is \[z^2\], if you square \[z^2\] you get \[(z^2)^2\]

OpenStudy (anonymous):

which equals \[z^4\], so your equation is \[(z^2 -7)(z^2 - 7)\]

OpenStudy (anonymous):

or \[(z^2 - 7)^2\]

OpenStudy (anonymous):

so when i solve these, i should solve the last term first?

OpenStudy (anonymous):

what do you mean "last term first"?

OpenStudy (anonymous):

solve for 49

OpenStudy (anonymous):

the last term is "c" if you remember the form of a quadratic equation: \[ax^2 +bx + c\]

OpenStudy (anonymous):

well if you remember, to solve these types of equations, you need to find two numbers whose product is "c" (+/- depending on the sign before it) and whose sum is b (also +/- depending on the sign before it)

OpenStudy (anonymous):

you don't really solve for the last term, you use it as a basis, since you'll only get half of the answer, basically

OpenStudy (anonymous):

and your variable would just be whatever the middle variable is (in ur case it was z^2) just make sure it is in quadratic form (that is, \[\sqrt{x^2} = x\]

OpenStudy (anonymous):

in ur case: \[\sqrt{z^4} = z^2\]

OpenStudy (anonymous):

and that holds true for the middle variable being your leading variable in the binomials in your answer.